# Strategic Content Creation with GenAI: To Share or Not to Share?

### v2

@@ 1.1 Our Contribution
Theorem 1 (Informal version of Theorem 8).
@@ 3.2 Existence of FSE
First, we consider the two extreme cases, $\rho=0$ and $\rho=1$. The former is trivial: creators can only be worse off from sharing; hence, withholding content is a dominant strategy, and no FSE exists in $G(0)$. The latter case is more interesting. If $\rho=1$, the platform allocates its entire GenAI-driven revenue to the creators. While this might not be practical, it provides a useful reference point for creator-centered ecosystems.
@@ 3.3 Analysis under Power Cost Functions
The proof of Proposition˜6 leverages previous work [50, Theorem 3] to construct a lower bound on $\left\|\mathcal{X}_{-i}(\rho)\right\|$ which depends on the number of creators $n$. We then obtain the desired result by substituting this lower bound in Theorem˜5 and extracting the necessary $n$.
@@ 4.2 Algorithm
We propose a simple search-based procedure for computing an approximate solution of PO, illustrated in Algorithm˜1. It receives as input an instance $\mathcal{I}$ and two parameters, $\delta$ and $\eta$. We carefully determine the values of $\delta$ and $\eta$ to reach a certain approximation level; we elaborate later in Theorem 8. Algorithm˜1 performs a grid search over the interval $[0,1]$, evaluating a discrete set of allocations parameters at resolution $\delta$. Starting with initialized variables (Line 3), the algorithm iterates through candidate values $\rho=\nicefrac{{t}}{{\left\lfloor 1/\delta\right\rfloor}}$ (Lines 48). For each candidate, it approximates the corresponding ESE (Line 6) and verifies whether it constitutes an $\eta$-FSE (Line 7). If this stability criterion is met, it updates the current best solution whenever the resulting platform revenue exceeds the previous maximum (Line 8). The algorithm ultimately returns the best-performing $\rho$ across all valid candidates (Line 9). The next result establishes that the algorithm returns a near-optimal solution with provable guarantees on both approximation quality and runtime.
@@ 4.3 Proof Sketch of Theorem˜8
Fix any instance $\mathcal{I}$ and $\varepsilon>0$, and let $\delta,\eta$ denote the parameters in Theorem˜8 that are passed to the algorithm. Next, focus on any feasible solution $\rho^{\star}$ to PO, and let $\rho$ denote the closest value to $\rho^{\star}$ the algorithm considers in Line 5. Notably, $\left|\rho-\rho^{\star}\right|\leq\delta$. Figure 1 illustrates how $\rho^{\star}$ is mapped to the unique ESE of $G(\rho^{\star})$, $\mathcal{X}(\rho^{\star})$. Similarly, $\rho$ is mapped to $\mathcal{X}(\rho)$, the unique ESE of $G(\rho)$.
@@ 6.4 GenAI-Dependent Traffic
$T(\mathbf{x},\mathbf{s})=\mu\left(\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})\right)^{\gamma}.$ (3)
@@ Results
Under this extended traffic model, the creator’s optimal sharing decision retains its threshold structure, with an adjusted threshold given by $\tau_{i}=\frac{1}{\alpha}\left[\left(\frac{\sum_{j}x_{j}+\alpha\sum_{j}x_{j}s_{j}}{\sum_{j}x_{j}+\alpha\sum_{j\neq i}x_{j}s_{j}}\right)^{1-\gamma}-1\right]$. Notably, larger values of $\gamma$ decrease $\tau_{i}$, making sharing more attractive. This occurs because higher traffic elasticity amplifies the positive externality that sharing exerts on the platform’s traffic.
@@ A.1 ESE Computation
Algorithm 2 provides a theoretically justified method for approximating the unique ESE for a given revenue-allocation parameter $\rho$. It employs multi-agent mirror descent with gradient-based updates to iteratively refine each creator’s quality choice $x_{i}$ while maintaining non-negativity constraints. The strong monotonicity of the enforced sharing game (established in Lemma˜3) guarantees convergence at a rate of $O(1/T)$ [12], where $T$ is the number of iterations.
@@
While Algorithm 2 provides provable convergence guarantees, we adopted a more practical approach in our simulations. Since each creator’s utility is strictly concave in their own quality $x_{i}$ (as established in Lemma˜3), the ESE is uniquely characterized by the first-order conditions: for each creator $i\in[n]$, the equilibrium quality satisfies
@@ A.2 FSE Stability Check
Algorithm 3 verifies whether a given profile is an FSE by checking if any creator can profitably deviate to not sharing ($s_{i}=0$). This is necessary and sufficient due to Lemma˜2: any profitable deviation is dominated by one with $s_{i}\in\{0,1\}$, and deviations with $s_{i}=1$ are already ruled out by the ESE property. For each creator, we solve a one-dimensional optimization problem to find their best response when withholding content. The profile is deemed stable only if no creator benefits from such deviation beyond the tolerance threshold $\varepsilon_{tol}$.
@@ A.3 Simulation Pipeline and Computational Resources
Appendix B Omitted Proofs from Section 3
This section is divided into two parts. In the first part, we state and prove all the auxiliary lemmas necessary for the proofs of the main results presented in Section 3. The second part provides the complete proofs of these main results.
@@ Appendix B Omitted Proofs from Section 3
B.1 Auxiliary Lemmas for Section 3’s Main Results
@@
By Proposition 2, the optimal sharing level for any quality $x_{i}^{\prime}$ must be either $s_{i}^{\prime}=0$ or $s_{i}^{\prime}=1$, with other values being weakly dominated. We analyze both cases and show that in each of them, the revenue is bounded above by $\mu(x_{i}^{\prime})^{\gamma}$. Thus, any deviation $(x_{i}^{\prime},s_{i}^{\prime})$ with $x_{i}^{\prime}>x_{i}^{\max}$ leads to negative utility, as for those values the cost $c_{i}(x_{i}^{\prime})$ exceeds the revenue.
@@
Let $j^{\star}=\text{arg}\min{j\in[n]}q_{j}$. Since $q_{j^{\star}}\leq\frac{1}{n}\leq\frac{1}{2}$, we have:
@@
Now, we divide the FOC of any creator $i$, given by Equation˜4, by the FOC of creator $j^{\star}$:
$\frac{1-(1-\gamma)q_{i}}{1-(1-\gamma)q_{j^{\star}}}=\frac{a_{i}\theta(\mathcal{X}_{i}(\rho))^{\theta-1}}{a_{j^{\star}}\theta(\mathcal{X}_{j^{\star}}(\rho))^{\theta-1}}$
@@
Notice that this is true for every creator $j$, not just $j^{\star}$. Substituting Equation˜5 back into the definition of $q_{i}$, we get:
@@
Substituting Equation˜7 back into the FOC of any creator $i$ given by Equation˜4, we have:
@@
Summing Equation˜8 over all creators $i\in[n]$, we get:
@@
Additionally, since $q_{i}\leq\frac{C}{C+1}<1$ for every creator $i$ by Equation˜6, it means that $\left\|\mathcal{X}_{-i}(\rho)\right\|=\Theta\left(\left\|\mathcal{X}(\rho)\right\|\right)=\Theta\left(n^{\frac{\theta-1}{\theta-\gamma}}\right)$ as well. This completes the proof of Lemma˜16. ∎
B.2 Proofs of Section 3’s Main Results
@@
$U_{i}(s_{i})$ $=T\cdot\frac{x_{i}}{D(s_{i})}+T\cdot\frac{\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})}{D(s_{i})}\cdot\rho\frac{x_{i}s_{i}}{x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}-c_{i}(x_{i}).$
@@
Notice that this simplified expression is consistent with Equation˜9 when $s_{i}=0$. Next, we compute the derivative of $U_{i}(s_{i})$ with respect to $s_{i}$:
@@
$=K\left\|\mathbf{x}\right\|^{\gamma-2}\left[\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|\right]$
@@
$=K(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-3}\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right]$
@@
By Proposition 2, for any fixed quality $x_{i}>0$, full sharing is the unique optimal choice if and only if:
@@
Appendix C Omitted Proofs From Section 4
@@ C.1 ESE Quality Lower Bound and Notations
$x_{i}^{\min}\coloneqq\left(c_{i}^{\prime}\right)^{-1}\left(\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\right).$
@@
$=\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\left[1+(\gamma-1)\frac{\mathcal{X}_{i}(\rho)}{\left\|\mathcal{X}(\rho)\right\|}\right]$
$\geq\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\gamma$
$\geq\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma,$
@@
$\mathcal{X}_{i}(\rho)\geq\left(c_{i}^{\prime}\right)^{-1}\left(\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\right)=x_{i}^{\min}.$
@@ C.2 Technical Lipschitz Continuity Results
$\left|\frac{\partial}{\partial z}\left(-A(y,z,\sigma)\right)\right|\leq\mu(Y_{LB})^{\gamma-2}(Y_{UB}+x^{\max})+M_{c^{\prime}}\leq B(1,1,0,1,2-\gamma)$
@@ Conclusion
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2,0,1,2-\gamma)\left|y_{1}-y_{2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|$
@@
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2,0,1,2-\gamma)\left|y_{1}-y_{2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|$
@@
According to Proposition 2, the optimal deviation involves either full sharing or no sharing. Thus, we can express the gain function as:
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq B(1,1,0,1,2-\gamma)\delta+B(0,2+\alpha,0,2,3-\gamma)\sqrt{n}\delta+B(0,2,0,1,1-\gamma)\delta\\$
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq\delta\cdot\left(B(1,1,0,1,2-\gamma)+B(0,(2+\alpha),0.5,2,3-\gamma)+B(0,2,0,1,1-\gamma)\right)$
@@ C.3 Lipschitz Continuity Results Used in the Proof of Algorithm˜1
Define $\bar{\mathbf{x}}\in\{\mathbf{x},\mathbf{x}^{\prime}\}$ to be the ESE profile among the two. Let $\phi_{k}(\mathbf{z},\sigma)$ denote the maximum utility gain player $k$ can achieve by unilaterally deviating from the profile $(\mathbf{z},\mathbf{1})$ when the platform’s revenue sharing parameter is $\sigma$. To apply Lemma 22, we need to establish bounds on other players’ contributions in the profiles $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
From Lemma 18, we know that $\left\|\bar{\mathbf{x}}_{-k}\right\|\geq\sum_{i}x_{i}^{\min}-\max_{k}x_{k}^{\min}\eqqcolon Y_{LB}$ for any player $k$ and any $\rho\in[0,1]$. Additionally, we know from Lemma 13 that $\left\|\bar{\mathbf{x}}_{-k}\right\|\leq\sum_{j=1}^{n}x_{j}^{\max}-\min_{k}x_{k}^{\max}\eqqcolon Y_{UB}$.
@@
Now, applying Lemma 22 with $\tilde{Y}_{LB},\tilde{Y}_{UB}$ as the bounds for other players’ contributions, we get:
@@
Without loss of generality, we assume $\mathbf{x}$ is the unique ESE in $G(\rho)$. Following the same steps as in the proof of Lemma˜24, we get that $0.5X_{LB}\leq\left\|\mathbf{x}\right\|\leq 2X_{UB}$, as well as $\left\|\mathbf{x}^{\prime}\right\|$ bounds. Next, we apply Lemma 23 with those bounds to get:

### v4

@@ 1.1 Our Contribution
Theorem 1 (Informal version of Theorem˜8).
@@ 3.2 Existence of FSE
First, we consider the two extreme cases, $\rho=0$ and $\rho=1$. The former is trivial—reators can only be worse off from sharing; hence, withholding content is a dominant strategy, and no FSE exists in $G(0)$. The latter case is more interesting. If $\rho=1$, the platform allocates its entire GenAI-driven revenue to the creators. While this might not be practical, it provides a useful reference point for creator-centered ecosystems.
@@ 3.3 Analysis under Power Cost Functions
The proof of Proposition˜6 involves the construction of a lower bound on $\left\|\mathcal{X}_{-i}(\rho)\right\|$ that grows with the number of creators $n$. We then obtain the desired result by substituting this lower bound in Theorem˜5 and extracting the necessary $n$.
@@ 4.2 Algorithm
We propose a simple search-based procedure for computing an approximate solution of PO, illustrated in Algorithm˜1. It receives as input an instance $\mathcal{I}$ and two parameters, $\delta$ and $\eta$. We carefully determine the values of $\delta$ and $\eta$ to reach a certain approximation level; we elaborate later in Theorem˜8. Algorithm˜1 performs a grid search over the interval $[0,1]$, evaluating a discrete set of allocations parameters at resolution $\delta$. Starting with initialized variables (Line 3), the algorithm iterates through candidate values $\rho=\nicefrac{{t}}{{\left\lfloor 1/\delta\right\rfloor}}$ (Lines 48). For each candidate, it approximates the corresponding ESE (Line 6) and verifies whether it constitutes an $\eta$-FSE (Line 7). If this stability criterion is met, it updates the current best solution whenever the resulting platform revenue exceeds the previous maximum (Line 8). The algorithm ultimately returns the best-performing $\rho$ across all valid candidates (Line 9). The next result establishes that the algorithm returns a near-optimal solution with provable guarantees on both approximation quality and runtime.
@@ 4.3 Proof Sketch of Theorem˜8
Fix any instance $\mathcal{I}$ and $\varepsilon>0$, and let $\delta,\eta$ denote the parameters in Theorem˜8 that are passed to the algorithm. Next, focus on any feasible solution $\rho^{\star}$ to PO, and let $\rho$ denote the closest value to $\rho^{\star}$ the algorithm considers in Line 5. Notably, $\left|\rho-\rho^{\star}\right|\leq\delta$. Figure˜1 illustrates how $\rho^{\star}$ is mapped to the unique ESE of $G(\rho^{\star})$, $\mathcal{X}(\rho^{\star})$. Similarly, $\rho$ is mapped to $\mathcal{X}(\rho)$, the unique ESE of $G(\rho)$.
@@ 6.4 GenAI-Dependent Traffic
$T(\mathbf{x},\mathbf{s})=\mu\left(\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})\right)^{\gamma}.$
@@ Results
Under this extended traffic model, the creator’s optimal sharing decision retains its threshold structure, with an adjusted threshold given by $\tau_{i}=\frac{1}{\alpha}\left[\left(\frac{\left\|\mathbf{x}\right\|+\alpha\left(x_{i}+\sum_{j\neq i}^{n}x_{j}s_{j}\right)}{\left\|\mathbf{x}\right\|+\alpha\sum_{j\neq i}x_{j}s_{j}}\right)^{1-\gamma}-1\right]$. Notably, larger values of $\gamma$ decrease $\tau_{i}$, making sharing more attractive. This occurs because higher traffic elasticity amplifies the positive externality that sharing exerts on the platform’s traffic.
@@ A.1 ESE Computation
Algorithm˜2 provides a theoretically justified method for approximating the unique ESE for a given revenue-allocation parameter $\rho$. It employs multi-agent mirror descent with gradient-based updates to iteratively refine each creator’s quality choice $x_{i}$ while maintaining non-negativity constraints. The strong monotonicity of the enforced sharing game (established in Lemma˜3) guarantees convergence at a rate of $O(1/T)$ [12], where $T$ is the number of iterations.
@@
While Algorithm˜2 provides provable convergence guarantees, we adopted a more practical approach in our simulations. Since each creator’s utility is strictly concave in their own quality $x_{i}$ (as established in Lemma˜3), the ESE is uniquely characterized by the first-order conditions: For each creator $i\in[n]$, the equilibrium quality satisfies
@@ A.2 FSE Stability Check
Algorithm˜3 verifies whether a given profile is an FSE by checking if any creator can profitably deviate to not sharing ($s_{i}=0$). This is necessary and sufficient due to Lemma˜2: any profitable deviation is dominated by one with $s_{i}\in\{0,1\}$, and deviations with $s_{i}=1$ are already ruled out by the ESE property. For each creator, we solve a one-dimensional optimization problem to find their best response when withholding content. The profile is deemed stable only if no creator benefits from such deviation beyond the tolerance threshold $\varepsilon_{tol}$.
@@ A.3 Simulation Pipeline and Computational Resources
Appendix B Omitted Proofs from Section˜3
This section is divided into two parts. In the first part, we state and prove all the auxiliary lemmas necessary for the proofs of the main results presented in Section˜3. The second part provides the complete proofs of these main results.
@@ Appendix B Omitted Proofs from Section˜3
B.1 Auxiliary Lemmas for Section˜3’s Main Results
@@
By Lemma˜2, the optimal sharing level for any quality $x_{i}^{\prime}$ must be either $s_{i}^{\prime}=0$ or $s_{i}^{\prime}=1$, with other values being weakly dominated. We analyze both cases and show that in each of them, the revenue is bounded above by $\mu(x_{i}^{\prime})^{\gamma}$. Thus, any deviation $(x_{i}^{\prime},s_{i}^{\prime})$ with $x_{i}^{\prime}>x_{i}^{\max}$ leads to negative utility, as for those values the cost $c_{i}(x_{i}^{\prime})$ exceeds the revenue.
@@
Let $j^{\star}=\arg\min_{j\in[n]}q_{j}$. Since $q_{j^{\star}}\leq\frac{1}{n}\leq\frac{1}{2}$, we have:
@@
Now, we divide the FOC of any creator $i$, given by Equation˜3, by the FOC of creator $j^{\star}$:
$\frac{1-(1-\gamma)q_{i}}{1-(1-\gamma)q_{j^{\star}}}=\frac{a_{i}\theta(\mathcal{X}_{i}(\rho))^{\theta-1}}{a_{j^{\star}}\theta(\mathcal{X}_{j^{\star}}(\rho))^{\theta-1}}.$
@@
Notice that this is true for every creator $j$, not just $j^{\star}$. Substituting Equation˜4 back into the definition of $q_{i}$, we get:
@@
Substituting Equation˜6 back into the FOC of any creator $i$ given by Equation˜3, we have:
@@
Summing Equation˜7 over all creators $i\in[n]$, we get:
@@
Additionally, since $q_{i}\leq\frac{C}{C+1}<1$ for every creator $i$ by Equation˜5, it means that $\left\|\mathcal{X}_{-i}(\rho)\right\|=\Theta\left(\left\|\mathcal{X}(\rho)\right\|\right)=\Theta\left(n^{\frac{\theta-1}{\theta-\gamma}}\right)$ as well. This completes the proof of Lemma˜16. ∎
B.2 Proofs of Section˜3’s Main Results
@@
$U_{i}(s_{i})$ $=T\cdot\frac{x_{i}}{D(s_{i})}+T\cdot\frac{\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})}{D(s_{i})}\cdot\rho\frac{x_{i}s_{i}}{x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}-c_{i}(x_{i})$
@@
Notice that this simplified expression is consistent with Equation˜8 when $s_{i}=0$. Next, we compute the derivative of $U_{i}(s_{i})$ with respect to $s_{i}$:
@@
$=K\left\|\mathbf{x}\right\|^{\gamma-2}\left[\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|\right].$
@@
$=K(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-3}\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right].$
@@
By Lemma˜2, for any fixed quality $x_{i}>0$, full sharing is the unique optimal choice if and only if:
@@
Appendix C Omitted Proofs From Section˜4
@@ C.1 ESE Quality Lower Bound and Notations
$x_{i}^{\min}\coloneqq\left(c_{i}^{\prime}\right)^{-1}\left(\frac{\mu}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\right).$
@@
$\geq\frac{\mu}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\left[1+(\gamma-1)\frac{\mathcal{X}_{i}(\rho)}{\left\|\mathcal{X}(\rho)\right\|}\right]$
$\geq\frac{\mu}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\gamma$
$\geq\frac{\mu}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma,$
@@
$\mathcal{X}_{i}(\rho)\geq\left(c_{i}^{\prime}\right)^{-1}\left(\frac{\mu}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\right)=x_{i}^{\min}.$
@@ C.2 Technical Lipschitz Continuity Results
$\left|\frac{\partial}{\partial z}\left(-A(y,z,\sigma)\right)\right|\leq\mu(Y_{LB})^{\gamma-2}(Y_{UB}+x^{\max})+M_{c^{\prime}}\leq B(1,1,0,1,2-\gamma).$
@@ Conclusion
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2,0,1,2-\gamma)\left|y_{1}-y_{2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|.$
@@
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2,0,1,2-\gamma)\left|y_{1}-y_{2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|.$
@@
According to Lemma˜2, the optimal deviation involves either full sharing or no sharing. Thus, we can express the gain function as:
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq B(1,1,0,1,2-\gamma)\delta+B(0,2+\alpha,0,2,3-\gamma)\sqrt{n}\delta+B(0,2,0,1,1-\gamma)\delta.$
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq\delta\cdot\left(B(1,1,0,1,2-\gamma)+B(0,(2+\alpha),0.5,2,3-\gamma)+B(0,2,0,1,1-\gamma)\right).$
@@ C.3 Lipschitz Continuity Results Used in the Proof of Algorithm˜1
Define $\bar{\mathbf{x}}\in\{\mathbf{x},\mathbf{x}^{\prime}\}$ to be the ESE profile among the two. Let $\phi_{k}(\mathbf{z},\sigma)$ denote the maximum utility gain player $k$ can achieve by unilaterally deviating from the profile $(\mathbf{z},\mathbf{1})$ when the platform’s revenue sharing parameter is $\sigma$. To apply Lemma˜22, we need to establish bounds on other players’ contributions in the profiles $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
From Lemma˜18, we know that $\left\|\bar{\mathbf{x}}_{-k}\right\|\geq\sum_{i}x_{i}^{\min}-\max_{k}x_{k}^{\min}\eqqcolon Y_{LB}$ for any player $k$ and any $\rho\in[0,1]$. Additionally, we know from Lemma˜13 that $\left\|\bar{\mathbf{x}}_{-k}\right\|\leq\sum_{j=1}^{n}x_{j}^{\max}-\min_{k}x_{k}^{\max}\eqqcolon Y_{UB}$.
@@
Now, applying Lemma˜22 with $\tilde{Y}_{LB},\tilde{Y}_{UB}$ as the bounds for other players’ contributions, we get:
@@
Without loss of generality, we assume $\mathbf{x}$ is the unique ESE in $G(\rho)$. Following the same steps as in the proof of Lemma˜24, we get that $0.5X_{LB}\leq\left\|\mathbf{x}\right\|\leq 2X_{UB}$, as well as $\left\|\mathbf{x}^{\prime}\right\|$ bounds. Next, we apply Lemma˜23 with those bounds to get: